A baseball player swings a bat. The bat is 1.5m from the fulcrum (axis of rotation) to the top of the bat. The start angle of the swing is 70° and the angle at which the ball is hit is 190°. The start time is 0s and the end time is 0.5s. The bat swings in the negative direction. What is the linear distance traveled by the end (top) of the bat?
Answer #1

The top of the bat follows a circular motion.

It is 1.5m from the center of the motion.

the motion makes an arc of final angle - initial angle = (- 170) - 70 = -240 degreess (Its start angle is 70 degrees and it goes in a negetive direction to finally rest at 190 or -170 degrees)

240 degrees is 3/4 of a full circle (240/360 = 3/4)

So, the linear distance travelled by top of the bat is 3/4 of a full circle

linear distance = 3/4 * 2*pi*r = 3/4 * 2 * 3.14 * 1.5 = 7.06m